OpenStudy (anonymous):

Use the Factor Theorem to determine whether x-2 is a factor of P(x)= x^4-2x^3+3x-6. Specifically, evaluate P at the proper value, and then determine whether x-2 is a factor.

jimthompson5910 (jim_thompson5910):

Hint: if you can show that P(2) = 0, then you will have shown that x - 2 is a factor

OpenStudy (anonymous):

Hm, I'm reading the example in the book, trying to see how to show that and uuhhhh math is not my strong point.

jimthompson5910 (jim_thompson5910):

P(x)= x^4-2x^3+3x-6 P(2)= (2)^4-2(2)^3+3(2)-6 P(2) = ??

OpenStudy (anonymous):

wait, i think i get it now...hold on

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

It's true p=0. You were telling me to plug 2 for x with the hint and i just didn't get it. Sometimes I have to see it...lol

OpenStudy (anonymous):

p(2)=2^4-2*2^3+2-6 2P=16-16+6-6 2p=0

jimthompson5910 (jim_thompson5910):

yes so because P(2) = 0, this means x-2 is a factor of P(x) by the factor theorem

jimthompson5910 (jim_thompson5910):

no you leave it as P(2) that's not the same as 2P

OpenStudy (anonymous):

Sorry, typing fast.

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome